# How General Are Measures of Choice Consistency? Evidence from Experimental and Scanner Data

### v1

@@ 5.4 Heuristic Rules
In addition to preferences, decision makers may use different heuristic rules in response to different environments. In the environment of budgetary choices, Choi et al. (2006) show that subjects’ portfolio choices can be explained by some simple rules, such as a diversification heuristic—allocating the same number of points to the two accounts. Halevy and Mayraz (2024) design simple investment rules for selecting portfolios and show that most of the subjects prefer to make allocations through the rule-based interface. It is possible that in the abstract environment of experiments, participants find the decisions to be difficult and adopt a set of heuristic rules to simplify their choices. In food purchasing decisions, consumers may use different heuristic rules such as habits formed in their daily life (Havranek et al., 2017), or reference dependence in perceiving price differences (Thaler and Shefrin, 1981).
@@
To explore such heuristic rules, in Experiment 2, we focus on “middle choosers” who choose near-middle options when the prices for two goods are similar. [^16] Specifically, we define middle choosers as participants who select the near-middle options—allocating between 40% to 60% budget to both goods (choosing the 5th, 6th or 7th option on the interface)—across six rounds in the risk task or three rounds in the social task where price ratios ($p_{1}/p_{2}$) range from 0.9 to 1.1. We cannot identify middle choosers in the food task as no price ratios fall within the range. The results reveal that middle choosers constitute 26.49% of participants in both risk and social tasks. Notably, conditional on being middle choosers in the risk task, the likelihood of being middle choosers in the social task is 45.00%, while this likelihood is only 19.82% among non-middle choosers in the risk task ($p<0.01$, two-sample proportion test).
@@
Another possible heuristic is “downward-sloping demand”, considering the degree of demand change in response to price change. Specifically, we calculate the Spearman’s rank correlation between $\ln(x_{1}/x_{2})$ and $\ln(p_{1}/p_{2})$ for each individual to measure the compliance with “downward-sloping demand” property (Echenique et al., 2023). The results show that pairwise correlations range from 0.324 to 0.539 in Experiment 2 ($p<0.01$ for all combinations) (See Panel B in Table 5). [^17] In addition, we also observe certain correlations of estimated parameters, $\hat{\alpha}$ and $\hat{\rho}$, among tasks within Experiment 2 (Table A14). This is notable because individuals do not necessarily respond consistently to prices due to specific utility functions in particular preference tasks. By contrast, “downward-sloping demand” measures have low or very low correlations in the scanner data.
@@
As the selection of middle options and the adherence to downward-sloping demand can be rationalized by certain utility functions, the observed correlations may be influenced by an underlying connection between risk preference and social preference. However, we believe that these correlations are more likely to arise from similar decision-making modes or heuristics used in response to the comparable budgetary interface presented in both risk and social tasks.
@@ 6 Concluding Remarks
Integrating both experimental data and scanner data, we conduct a comprehensive study to investigate the generalizability of consistency scores. Our primary findings reveal a striking lack of correlation between individuals’ consistency scores for risky decisions made in the experiment and food consumption choices observed in supermarkets. Within the supermarket context, we observe generally low correlations among food consumption decisions across categories and years; within the experimental setting, we observe on average moderate correlations among different tasks using budgetary interface. We further show that these correlational patterns remain robust when using a wide range of indices of consistency proposed in the literature. Our study adds to the understanding that preferences and consistency derived from observable choices do not necessarily generalize across different environments.
@@
We delve into several unobservable factors that may underpin the consistency scores across different environments, including randomness in choice behaviors, budget constraints, formation of preferences, and heuristic rules. In line with this, Choi et al. (2014) highlight the complexities of measuring decision quality, and argue that assessing decision quality in real-world contexts is challenging because decision makers “might have different preferences over the same outcomes, or face different but unobserved incentives and constraints, or have different information, or hold different beliefs.” To address these issues, they propose the use of choice consistency as a measure of decision-making quality, particularly in relation to departures from GARP using experimental data. Their study, along with several subsequent ones, further demonstrate that consistency measure holds real-world significance, even after accounting for unobserved constraints, preferences, and beliefs. Here we present additional results to pin down some underlying factors of consistency measures. In experimental data, we find that consistency is shaped by learning effects, heuristic rules of choosing the middle option and responding to price changes, and the cognitive ability of individuals. In the scanner data, consistency is influenced by contextual factors such as seasonality, the time of day, promotional discounts, as well as individual shopping experiences.
@@
Overall, these findings suggest that the observed lack of correlation in consistency scores between the two settings may stem partly from measurement errors, particularly in how budget constraints are constructed, as well as the influence of contextual factors. This disconnect may be partly due to the multidimensional nature of decision-making quality as well. Specifically, participants in the experiment are often faced with novel and abstract problems, while consumers in supermarkets typically make repeated daily decisions in familiar environments. This distinction closely resembles the difference between the decision-making capacity of a routine task and a problem-solving task (Deming, 2021). It is also reflected in the contrast between fluid intelligence, which is the ability to reason and solve problems in unfamiliar situations, and crystallized intelligence, which involves accumulated knowledge, skills, and experiences (Cattell, 1943).
@@
To sum, our study contributes to a fundamental aspect of economics—choice consistency within the framework of utility maximization under budget constraints (Samuelson, 1938; Afriat, 1967; Varian, 1982). It highlights a disconnect between experimental settings and supermarket environments, moderate correlations within experimental tasks and weak to moderate correlations in supermarket settings. Our results suggest that revealed preference measures have yet to capture the intricate and multidimensional nature in decision-making. Looking ahead, it is crucial for future research to explore alternative measures and frameworks of choice consistency that can more accurately capture the complexities of decision making in controlled laboratory experiments, field experiments, and everyday life.
@@ Appendix A Additional Figures and Tables
Figure A1: CCEI Scores Across Tasks in Experiment 2
Figure A2: CCEI Scores in Scanner Dataset 2
@@
Family Monthly Income $<$ RMB 5,000 24.4% 19.2%
@@
$>$ RMB 10,000 35.6% 51.0%
Education $<$ High school 44.9% 36.8%
@@
Exp: Risk Scanner: Food Correlation
Panel A: Baseline indices
@@
Panel B: Heterogeneous power for budget constraints
@@
Panel A: Baseline indices
@@
LL 0.522*** 0.379*** 0.383***
  • Notes: $N=302$. Spearman’s rank correlations are reported. Each coefficient is calculated using the index as a measure for both consistency scores. Two exceptions are FOSD and EU, which are applied exclusively to the risk task; for these, we calculate the correlations between them and CCEI scores for other tasks. -- denotes that the measure is not applicable. * $p<0.10$, ** $p<0.05$, *** $p<0.01$.
@@
Panel A: Baseline indices
@@
Panel B: Heterogeneous power for budget constraints
@@
Table A8: Correlations of Alternative Indices for Budgets Constraints and Heuristic Rules
@@
  • Notes: Spearman’s rank correlations are reported. -- denotes that the measure is not applicable. * $p<0.10$, ** $p<0.05$, *** $p<0.01$.
@@
Table A10: Pairwise Comparisons Across Difference Combination of Seasons in Scanner Dataset 1
@@
Season: $s_{1}$ Season: $s_{2}$ Mean ($s_{1}$) Mean ($s_{2}$) $p$-value Mean ($s_{1}$) Mean ($s_{2}$) $p$-value
Spring Summer 0.371 0.422 $0$6.356 -5.816 0.108
Spring Fall 0.371 0.445 $0$6.356 -6.303 0.886
Spring Winter 0.371 0.425 $0$6.356 -6.786 0.224

### v2

@@ 5.4 Heuristic Rules
In addition to preferences, decision makers may use different heuristic rules in response to different environments. In the environment of budgetary choices, Choi et al. (2006) show that the portfolio choices of subjects can be explained by some simple rules, such as a diversification heuristic, i.e., allocating the same number of points to the two accounts. Halevy and Mayraz (2024) design simple investment rules to select portfolios and show that most subjects prefer to make allocations through the rule-based interface. It is possible that in the abstract environment of experiments, participants find the decisions to be difficult and adopt a set of heuristic rules to simplify their choices. In food purchase decisions, consumers may use different heuristic rules, such as habits formed in their daily life (Havranek et al., 2017), or reference dependence to perceive price differences (Thaler and Shefrin, 1981).
@@
To explore such heuristic rules, in Experiment 2, we focus on “middle choosers” who choose near-middle options when the prices for two goods are similar. [^12] Specifically, we define middle choosers as participants who select the near-middle options—allocating between 40% to 60% budget to both goods (choosing the 5th, 6th or 7th option on the interface)—across six rounds in the risk task or three rounds in the social task where price ratios ($p_{1}/p_{2}$) range from 0.9 to 1.1. We cannot identify middle choosers in the food task as no price ratios fall within the range. The results reveal that middle choosers constitute 26.49% of participants in both risk and social tasks. Notably, conditional on being middle choosers in the risk task, the likelihood of being middle choosers in the social task is 45.00%, while this likelihood is only 19.82% among nonmiddle choosers in the risk task ($p<0.01$, two-sample proportion test).
@@
Another possible heuristic is “downward-sloping demand”, considering the degree of demand change in response to price change. Specifically, we calculate the Spearman’s rank correlation between $\ln(x_{1}/x_{2})$ and $\ln(p_{1}/p_{2})$ for each individual to measure compliance with “downward-sloping demand” property (Echenique et al., 2023). The results show that pairwise correlations range from 0.324 to 0.539 in Experiment 2 ($p<0.01$ for all combinations) (See Panel B in Table 5). [^13] In addition, we also observe certain correlations of estimated parameters, $\hat{\alpha}$ and $\hat{\rho}$, across tasks within Experiment 2 (Table A14). This is notable because individuals do not necessarily respond consistently to prices due to specific utility functions in particular preference tasks. In contrast, “downward-sloping demand” measures have low or very low correlations in the scanner data.
@@
As the selection of middle options and the adherence to downward-sloping demand can be rationalized by certain utility functions, the observed correlations may be influenced by an underlying connection among risk, social, and food preferences. However, we believe that these correlations are more likely to arise from similar decision-making modes or heuristics used in response to the comparable budgetary interface presented in risk, social, and food tasks.
5.5 Summary with a Conceptual Framework
To help understand our findings, we consider a conceptual framework as follows. The measured choice consistency $e$ is determined by the true unobserved choice consistency $e^{*}$, the preference factor $\theta$, and the measurement error $\nu$, which is due to measurement. Accordingly, we have
$e_{ij}=f(e^{*}_{ij},\theta_{ij},\nu_{ij}),$ (3)
where $i=1$ refers to the experimental setting, $i=2$ refers to the supermarket setting, and $j$ denotes the preference domain.
Measurement error is likely to be small in the experimental environment as evidenced by the small number of random DMs shown in Table 4. We observe in Result 2 that there are moderate correlations of observed choice consistency between different preference domains. Since risk, social and food preferences are presumably independent, the observed correlations between $e_{1j}$ are more likely to be driven by the underlying correlations of $e^{*}_{1j}$ between domains. That is, there is a common underlying choice consistency across preferences measured in the experimental setting.
Conversely, in the supermarket context, measurement error is likely to be substantial, given the concerns regarding the construction of budget lines discussed in Section 5.2 and the setting-specific factors outlined in Table 6. Nevertheless, the correlations between category combinations remain significant, albeit small (Result 3), and increase to 0.5 when more realistic preference restrictions are applied (Tables 3 and A7). These observations indicate that while measurement error is a non-negligible issue in the scanner data, it is not large enough to completely obscure the correlations of actual choice consistency. Therefore, measurement error alone cannot fully account for the lack of correlation between the experimental and scanner data. Collectively, these findings suggest that the underlying choice consistency is likely uncorrelated between settings, highlighting the multidimensional nature of choice consistency across different domains.
6 Predictive Validity
Our observations are in line with the notion that decision making capacity is multidimensional. To further substantiate our interpretation and to rule out measurement error as a full explanation, we examine the predictive validity of choice consistency in the two settings. Namely, we go beyond their specific measurement domains, and test whether these measures independently predict different aspects of consumer behavior. As Snowberg and Yariv (2025) note, “a measure has high predictive validity if it correlates with other variables that are known or expected to be theoretically related to it”. The degree of predictive validity can help shed light on whether the measure reflects meaningful variations in behavior rather than random or idiosyncratic noise.
In budgetary experiments, choice consistency reflects individuals’ capacity to solve novel problems and identify relevant trade-offs, and it is theoretically and empirically linked to financial consequences such as wealth accumulation (Choi et al., 2014; Carvalho et al., 2024). We therefore investigate whether consistency in the experiment predicts money-saving behavior. Money-saving behavior is measured by the frequency and magnitude of discount usage, including proportion of transactions with discounts, aggregate discount rate, and average transaction-level discount rate.
In the daily supermarket shopping, by contrast, choice consistency indicates the ability to form habits and to maintain routines across time periods in familiar environments (Cherchye et al., 2020; Crawford, 2010). Consequently, we examine how consistency in the supermarket relates to consumption regularity. Consumption regularity is measured by month-to-month volatility of shopping patterns across hours of day, days of week and ten-day periods of the month. In particular, for each individual in each year, let $\pi_{g}^{t}=\frac{v_{g}^{t}}{\sum_{g\in G}v_{g}^{t}}$, where $g$ denotes the group in the set $G$, $t$ denotes the month, and $v$ represents transaction amount or count. We consider three sets of groups separately defined along the temporal dimensions: hours of day, days of week, and ten-day periods of the month. Within group set $G$, we first calculate the standard deviation of each group $g$’s shares over 12 months, $s_{g}=\operatorname{sd}(\left\{\pi_{g}^{t}\right\}_{t=1}^{12})$, then average these standard deviations across all groups: $V_{G}=\frac{1}{|G|}\sum_{g\in G}s_{g}$. Higher $V_{G}$ indicates greater instability.
We link the consistency scores from Experiment 1 and Scanner Dataset 1 to consumers’ supermarket scanner records covering all product categories both within-sample (2019–2020) and out-of-sample (2021). We consider consumers’ in‑store purchases across all product categories. To mitigate the influence of outliers, we further exclude the top 0.5% transactions according to the transaction amount prior to discounts. In all regression analyses, we control for demographics, IQ, and Big Five personality traits.
We find a clear dissociation in Table 8. First, consistency in the experiment predicts money-saving behavior (Panel A). Namely, individuals with higher consistency scores with FOSD are significantly more likely to use discounts. [^14] Statistically significant effects for experimental measures are found with stricter criteria such as consistency with FOSD or EU maximization; the GARP alone shows no significant effect, which supports our mechanisms above: some individuals may use simple heuristics that satisfy GARP in the experiments but do not necessarily indicate high quality. This result is economically meaningful, robust across multiple discount metrics, and holds both within-sample (2019–2020) and out-of-sample (2021). In particular, a one-standard-deviation increase in FOSD is associated with a 7.1–7.8% standard-deviation increase in the proportion of transactions with discounts, a 3.6–7.2% standard deviation increase in aggregate discount rates, and a 4.9–7.4% standard-deviation increase in average transaction-level discount rate. In contrast, supermarket-based consistency measures rarely show a relationship with discount usage.
Table 8: OLS Regressions for Money-saving Behavior and Consumption Regularity
Panel A: Money-saving behavior Proportion of transactions with discounts Aggregate discount rate Average transaction-level discount rate 2019 2020 2021 2019 2020 2021 2019 2020 2021 (1) (2) (3) (4) (5) (6) (7) (8) (9) Exp: FOSD 0.045** 0.043*** 0.053*** 0.015 0.027** 0.020* 0.019* 0.021** 0.029*** (0.019) (0.017) (0.020) (0.012) (0.011) (0.011) (0.011) (0.011) (0.011) Scanner: CCEI -0.062 -0.045 -0.046 -0.072* -0.049 0.020 -0.048 -0.026 -0.040 (0.066) (0.059) (0.069) (0.039) (0.035) (0.036) (0.037) (0.035) (0.040) Constant 0.462*** 0.435*** 0.439*** 0.202*** 0.163*** 0.115*** 0.123*** 0.106*** 0.127*** (0.072) (0.064) (0.076) (0.044) (0.038) (0.041) (0.041) (0.038) (0.044) Controls Y Y Y Y Y Y Y Y Y N 1,052 1,046 1,038 1,052 1,046 1,038 1,052 1,046 1,038 R2 0.101 0.069 0.051 0.112 0.064 0.041 0.121 0.068 0.054 Panel B: Consumption regularity Hours of day Days of week Ten-day periods of the month 2019 2020 2021 2019 2020 2021 2019 2020 2021 (1) (2) (3) (4) (5) (6) (7) (8) (9) Exp: FOSD 0.003 0.001 0.000 0.008 -0.001 -0.013 0.005 0.018 -0.000 (0.003) (0.004) (0.004) (0.007) (0.009) (0.009) (0.010) (0.015) (0.017) Scanner: CCEI -0.043*** -0.039*** -0.035*** -0.072*** -0.075** -0.077*** -0.124*** -0.116** -0.101** (0.010) (0.012) (0.012) (0.023) (0.029) (0.027) (0.036) (0.049) (0.050) Constant 0.115*** 0.118*** 0.120*** 0.212*** 0.241*** 0.265*** 0.329*** 0.333*** 0.385*** (0.012) (0.013) (0.014) (0.025) (0.032) (0.030) (0.040) (0.054) (0.054) Controls Y Y Y Y Y Y Y Y Y N 1,050 1,043 1,017 1,050 1,043 1,017 1,050 1,043 1,017 R2 0.072 0.081 0.058 0.049 0.049 0.049 0.063 0.044 0.047 Notes: In the table, Panel A presents results for money-saving behaviors, with dependent variables being the proportion of transactions with discounts (Columns 1–3), aggregate discount rate (Columns 4–6), and average transaction-level discount rate (Columns 7–9). Panel B examines consumption regularity, with dependent variables measuring month-to-month volatility in shopping patterns across hours of day (Columns 1–3), days of week (Columns 4–6), and ten-day periods of the month (Columns 7–9). Higher (lower) volatility indicates lower (higher) consumption regularity. Controls includes IQ, Big Five personality traits, female, family size, dummies for medium and high income, dummies for medium and high education, and dummies for medium and elder age. Tables A15 and A16 report the results that include the coefficients of control variables. Robust standard errors are in parentheses. * $p<0.10$, ** $p<0.05$, *** $p<0.01$.
Conversely, consistency measure from the supermarket predicts consumption regularity (Table 8, Panel B). Consumers with higher consistency scores in the scanner data exhibit shopping routines with significantly less fluctuations, in both 2019–2020 and 2021. Their transaction amount in the supermarket is more stable in terms of hours and days across months, showing significantly less month-to-month volatility. Specifically, a one-standard-deviation increase in CCEI scores is associated with an 8.3–12.7% standard-deviation reduction in the month-to-month volatility measured using hours of day, a 7.7–9.5% standard-deviation reduction using days of week, and a 6.0–10.6% standard-deviation reduction using ten-day periods of the month. The results remain robust when using transaction counts instead of transaction amount to measure volatility (Table A17). Notably, none of the consistency measures from the experiment show significant correlations with any of these indicators of consumption regularity.
Overall, consistency in both the experimental and scanner data demonstrates predictive validity and correlates with distinct, economically meaningful consumer behaviors. This finding alleviates concerns that these measures may simply capture noise. Instead, it provides evidence in support of the multidimensional nature of choice consistency.
@@ 7 Concluding Remarks
Economic analysis posits that decision makers aim to maximize their utility within budget constraints, and revealed preference analysis provides a framework to evaluate the consistency in choice data. Moreover, consistency score is proposed as a measure of decision quality and is linked to wealth disparities, development gaps, and policy effectiveness. Building on this research, we conduct a comprehensive study to explore the generalizability of consistency scores by integrating both experimental data and scanner data. Our primary findings reveal a lack of correlation between individuals’ consistency scores for risky decisions made in the experiment and food consumption choices observed in supermarkets. Within the experimental setting, we observe on average moderate correlations among different tasks using budgetary interface; similarly, in the supermarket context, food consumption decisions across categories and years also show significantly positive correlations. We further show that these correlational patterns remain robust in both settings when using a wide range of consistency indices proposed in the literature. Our study adds to the understanding that preferences and consistency derived from observable choices do not necessarily generalize across different environments.
@@
We explore several unobservable factors that may underpin consistency scores in different environments, including randomness in choice behaviors, budget constraints, formation of preferences, and heuristic rules. In line with this, Choi et al. (2014) highlight the complexities of measuring decision quality and argue that assessing decision quality in real world contexts is challenging because decision makers “might have different preferences over the same outcomes, or face different but unobserved incentives and constraints, or have different information or hold different beliefs”. To address these issues, they propose the use of choice consistency as a measure of decision-making quality, particularly in relation to departures from GARP using experimental data. Their study, along with several subsequent ones, further demonstrates that the consistency measure holds real-world significance even after accounting for unobserved constraints, preferences, and beliefs. Here, we present additional results to pin down some underlying factors of consistency measures. In experimental data, we find that consistency is shaped by learning effects, heuristic rules of choosing the middle option and responding to price changes, and the cognitive ability of individuals. In scanner data, consistency is influenced by contextual factors such as seasonality, time of day, promotional discounts, as well as individual shopping experiences.
@@
In general, these findings suggest that the observed lack of correlation in consistency scores between the two settings may stem in part from measurement errors, particularly in how budget constraints are constructed, as well as the influence of contextual factors. This disconnect may also be due in part to the multidimensional nature of the quality of decision-making. Specifically, participants in the experiment are often faced with novel and abstract problems, while consumers in supermarkets typically make repeated daily decisions in familiar environments. To support this interpretation, we present additional evidence that consistency in both the experimental and scanner data effectively predicts distinct consumer behaviors. This distinction closely resembles the difference between the decision-making capacity of a routine task and a problem-solving task (Deming, 2021; Cerigioni, 2021; Ilut and Valchev, 2023). It is also reflected in the contrast between fluid intelligence, which is the ability to reason and solve problems in unfamiliar situations, and crystallized intelligence, which involves accumulated knowledge, skills, and experiences (Cattell, 1943).
@@
In summary, our study contributes to a fundamental aspect of economics: choice consistency within the framework of utility maximization under budget constraints (Samuelson, 1938; Afriat, 1967; Varian, 1982). It highlights a disconnect between experimental settings and supermarket environments, moderate correlations within experimental tasks, and weak to moderate correlations in supermarket settings. Our results suggest that the revealed preference measures have yet to capture the intricate and multidimensional nature of decision-making. Looking ahead, future research may explore alternative measures and frameworks of choice consistency that can more accurately capture the complexities of decision-making in controlled laboratory experiments, field experiments, and everyday life.
@@ Appendix A Additional Figures and Tables
Figure A1: Distribution of CCEI Scores in Experiment 2
Figure A2: Distribution of CCEI Scores in Scanner Dataset 2
@@
Family Monthly Income << RMB 5,000 24.4% 19.2%
@@
>> RMB 10,000 35.6% 51.0%
Education << High school 44.9% 36.8%
@@
Exp: Risk Scanner: Consumption Correlation
Panel A: Consistency indices
@@
Panel B: Power of the test
@@
Panel A: Consistency indices
@@
$L$ 0.522*** 0.379*** 0.383***
  • Notes: $N=302$. Spearman’s rank correlations are reported. Each coefficient is calculated using the index as a measure for both consistency scores. Two exceptions are FOSD and EU, which are applied exclusively to the risk task; for these, we calculate the correlations between them and CCEI scores for other tasks. - denotes that the measure is not applicable. * $p<0.10$, ** $p<0.05$, *** $p<0.01$.
@@
Panel A: Consistency indices
@@
Panel B: Power of the test
@@
Table A8: Correlations of Alternative Indices for Budget Constraints and Heuristic Rules
@@
  • Notes: Spearman’s rank correlations are reported. - denotes that the measure is not applicable. * $p<0.10$, ** $p<0.05$, *** $p<0.01$.
@@
Table A10: Pairwise Comparisons Across Different Combinations of Seasons in Scanner Dataset 1
@@
Season: $s_{1}$ Season: $s_{2}$ Mean: $s_{1}$ Mean: $s_{2}$ $p$-value Mean: $s_{1}$ Mean: $s_{2}$ $p$-value
Spring Summer 0.371 0.422 0.000 -6.356 -5.816 0.108
Spring Fall 0.371 0.445 0.000 -6.356 -6.303 0.886
Spring Winter 0.371 0.425 0.000 -6.356 -6.786 0.224